OpenStudy (anonymous):

Inverse Variation: Y=-12 X=5 Formula Set-Up( I think): 5(-12)=K I assume I'd be dividing them, so they'd turn out like this; y/x=-2.4 unless they are multiplying it'd be xy=-60. Am I right on how to solve this, by dividing? On another note; Assume that y varies inversely with/as x. If y = -3 when x = 2.4, what is y when x = 2<---When a problem is like that, how do I start it off?

OpenStudy (saifoo.khan):

y = k/x

OpenStudy (anonymous):

That is the formula for the second problem?

OpenStudy (saifoo.khan):

For both. since they both are inversely propotional.

OpenStudy (anonymous):

Okay, so is my first answer correct?

OpenStudy (saifoo.khan):

this is wrong, I assume I'd be dividing them, so they'd turn out like this; y/x=-2.4 unless they are multiplying it'd be xy=-60.

OpenStudy (anonymous):

Please explain, then.

OpenStudy (saifoo.khan):

The second step you did is perfect. k = -60 Now, y = k/x y = -60/x This is correct. But i dont get " y/x=-2.4 "

OpenStudy (anonymous):

Well, I was thinking at first the y/x would represent the original y and x which after dividing, came out to be -2.4. But, now I get how to do the problem after I turn k/5 into a fraction, factor it out by multiplying 5 and -12(5) = -60 right?

OpenStudy (saifoo.khan):

-60 thing is right.

OpenStudy (anonymous):

Thanks!

OpenStudy (saifoo.khan):

Welks!